Circular No. 8260

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003lb, 2003lc, AND 2003ld Additional apparent supernovae have been reported from CCD images:

SN 2003lb in UGC 2850 by M. Ory (Delemont, Switzerland, 0.61-m reflector), and SN 2003lc in UGC 934 and SN 2003ld in UGC 148 by T. Puckett and L. Cox (cf. IAUC 8225; 0.50-m reflector).

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003lb Dec. 25.817 3^{h}45^{m}05.^{s}67 +20^{o}45′39′′.3 15.8 8′′.6 E, 3′′.9 S
2003lc Dec. 26.03 1 23 25.82 +30 47 14.9 17.0 32′′.1 W, 10′′.7 N
2003ld Dec. 28.02 0 15 51.85 +16 05 21.6 16.8 8′′.6 E, 1′′.3 S

Ory notes that SN 2003lb was at mag R = 16.1 on an image taken on Dec. 26.767 UT, adding that nothing is visible at the location of SN 2003lb on a red Palomar Sky Survey plate from 1951 Dec. 1 or a U.K. Schmidt Telescope plate from 1991 Oct. 13. SNe 2003lc and 2003ld were confirmed on CCD frames taken by Puckett on Dec. 27.15 and 29.15, respectively, but both new objects were absent from his images taken during 1998–2002.

C. L. Gerardy, University of Texas at Austin, reports that a lowresolution optical spectrogram (resolution 600; range 420–950 nm) of SN 2003lb, obtained on Dec. 27.23 UT with the 2.7-m Harlan J. Smith Telescope (+ LCS spectrograph) at McDonald Observatory, shows it to be a type-Ia supernova near maximum light; adopting the 5425 km/s redshift of UGC 2850 from the NASA/IPAC Extragalactic Database (NED), the expansion velocity of the Si II feature (rest wavelength 635.5 nm) is 12400 km/s. M. Salvo, Australian National University and Mt. Stromlo Observatory, also reports that a spectrogram (useful range 370–610 nm) of SN 2003lb, taken in rather poor observing conditions with the 2.3-m telescope (+ Double Beam Spectrograph) on Dec. 27.5, shows it to be a type-Ia supernova around maximum light.

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2003lc, obtained by P. Berlind on Dec. 28.12 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum. Adopting the NED recession velocity of 10494 km/s for the host galaxy, the supernova expansion velocity is 10100 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova is 4 ± 2 days after maximum light.

2003 December 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green